# Planting and MCMC Sampling from the Potts model

### v1

@@ Appendix C Remaining Proofs for Section 4
Write $T=\{i_{1},\dots,i_{k}\}$, and let $1\leq j\leq k$. Suppose we conditioned on the “outcome” for $R$-neighbourhoods of $i_{1},\dots,i_{j-1}$: that is, the number of surviving leaves, and which of the vertices in $\bigcup_{\ell<j}V_{i_{\ell}}$ belong to the largest component.
Regardless of outcome on the previous neighbourhoods, there are at least $\chi n^{\prime}/2-d^{R}|S^{\prime}|\geq\chi n/3$ undetermined vertices in the largest component, and these are chosen, by symmetry, uniformly at random from the set of at most $n$ not yet determined vertices.
Now consider the number of surviving leaves, $L_{i_{j}}$, in the $R$-neighbourhood of $i_{j}$. By Corollary D.4, the probability that $L_{i_{j}}<Kd$ is $(\varphi p+1-p)^{d_{i_{j}}}+o_{R}(1)$ (as $(\varphi p+1-p)^{d_{i_{j}}})$ is the survival probability for the corresponding branching process). Taking $R$ large enough, we can get $o_{R}(1)\leq\epsilon_{1}/2$.
Then, conditional on $L_{i_{j}}\geq Kd$, $|V_{i_{j}}|\geq K(d-1)$, and the probability that neither of these $K(d-1)$ belong to a u.a.r. subset of $<n$ vertices of size $\geq\chi n/3$ is at most $(1-\chi/3)^{K(d-1)}\rightarrow 0$ as we increase $K$. Thus for $K$ large enough, this probability is at most $\epsilon_{1}/2$.
@@
We showed $\mathbb{P}(|T|<(1-\epsilon_{0})|S^{\prime}|)\leq n^{-C|S^{\prime}|}$, and $\mathbb{P}({\mathcal{C}}_{good}\mid|T|\geq(1-\epsilon_{0})|S^{\prime}|)\leq\mathrm{e}^{-\Omega(n)}$. (D.1) gives $\mathbb{P}(S^{\prime}\cap{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})=\emptyset\mid|T|\geq(1-\epsilon_{0})|S^{\prime}|,T)\leq\prod_{i\in T}[(\varphi p+1-p)^{d_{i}}+\epsilon_{1}]$. Remains to show that this is bounded by a sufficiently small probability depending on $S^{\prime}$, not $T$, for all $T\subseteq S^{\prime}$ with $|T|\geq(1-\epsilon_{0})|S^{\prime}|$, provided that we make $\epsilon_{0}$ and $\epsilon_{1}$ small enough. In particular, we will get that (for $\epsilon_{0},\epsilon_{1}$ sufficiently small), for all $T\subseteq S^{\prime}$ with $|T|\geq(1-\epsilon_{0})|S^{\prime}|$,
@@
As a first step towards (D.2), note that for each $i\in S^{\prime}$, $d_{i}\leq d$, and since also $(\varphi p+1-p)<1$, it follows that $(\varphi p+1-p)^{d_{i}}+\epsilon_{1}\leq(\varphi p+1-p)^{d_{i}}(1+\epsilon_{1}/(\varphi p+1-p)^{d})$.
Also, we can bound, for all sufficiently large $T$,
@@
Combining, we get
$\prod_{i\in T}[(\varphi p+1-p)^{d_{i}}+\epsilon_{1}]\leq[(1+\epsilon_{1}/(\varphi p+1-p)^{d})(\varphi p+1-p)^{d\epsilon_{0}}]^{|S^{\prime}|}\prod_{i\in S^{\prime}}(\varphi p+1-p)^{d_{i}}$
Note that by making $\epsilon_{0},\epsilon_{1}\rightarrow 0$, we get $(1+\frac{\epsilon_{1}}{(\varphi p+1-p)^{d}})(\varphi p+1-p)^{d\epsilon_{0}}\rightarrow 1$, so in particular we can have the $\epsilon_{0},\epsilon_{1}>0$ small enough so the expression becomes at most $21+\epsilon/$ 2, and hence we get (D.2).
@@
$[\epsilon(\varphi p+1-p)^{d}/2]^{|S^{\prime}|}+(1+\epsilon/2)^{|S^{\prime}|}\prod_{i\in S^{\prime}}(\varphi p+1-p)^{d_{i}}\leq(1+\epsilon)^{|S^{\prime}|}\prod_{i\in S^{\prime}}(\varphi p+1-p)^{d_{i}},$
@@
Appendix E Proof of Theorems 3.4 and 3.5
@@
In this section, we prove Theorems 3.4 and 3.5. The proofs of these follow closely those in [RCrandom] (which are in turn similar to previous arguments in [GS, BG]), so we outline the key ideas and only do a bit more carefully the probability amplification (which is not included therein).
@@
Also, for $\ell$ as above, there exists a constant $L>0$ such that for every vertex $v$, $B_{\ell}(v)$ is $L$-treelike, see, e.g., [BG, Fact 2.3.]. Consider any graph $G$ that satisfies these properties.
We next outline the theorem for $\varepsilon=\tfrac{1}{4}$, the full details for this part can be found in [RCrandom]. Then, we prove that probability amplification applies for any $\varepsilon=\mathrm{e}^{-\Omega(n)}$. Consider two copies of the random-cluster dynamics: (1) $(X_{t})_{t\geq 0}$ starting from all-out, and (2) $(\hat{X}_{t})_{t\geq 0}$ starting from $\pi_{G,\mathrm{d}}$ which is restricted to the disordered phase by ignoring updates that would make the configuration not disordered. Note that the stationary distribution of $\hat{X}_{t}$ is $\pi_{G,\mathrm{d}}$, and thus for all $t\geq 0$, $\hat{X}_{t}\sim\pi_{G,\mathrm{d}}$.
@@
$\mathbb{P}(X_{T}\neq\hat{X}_{T})\leq$ $m\mathbb{P}(\overline{{\mathcal{E}}_{<T}})+\sum_{e\in E(G)}\mathbb{P}(\mathds{1}\{e\in X_{T}\}\neq\mathds{1}\{e\in\hat{X}_{T}\}\mid{\mathcal{E}}_{<T}).$ (E.1)
@@
where $\pi^{-}_{G,v,\ell}$ is the distribution on $B_{\ell}(v)$ conditioned on an all-out boundary. Since $\mathbb{P}(e\in\hat{X}_{T})=\pi_{G,\mathrm{d}}(1_{e})$, WSM within the disordered phase bounds the third term by $\frac{1}{100m}$. The second term is also bounded by $\frac{1}{100m}$ using mixing time estimates, in this case log-Sobolev inequalities. We don’t give the full details here (which, as mentioned before, can also be found in [RCrandom, GS, BG]) but the idea is that, since the neighbourhood of $v$ is treelike (up to removing a constant number of edges), the log-Sobolev constant for $\pi_{G,v,\ell}^{-}(1_{e})$ is a constant factor away from that on the regular tree. This in turn is just percolation so the log-Sobolev constant for the chain $X_{v}^{t}$ is at most $C_{\mathrm{d}}/|V(B_{\ell}(v))|$ (in the ordered regime, the bound for the log-Sobolev constant on the tree with + boundary condition is given in [treeRC]). Since $\pi_{G,v,\ell}^{-}$ is the stationary distribution of the chain $X_{v}^{t}$, the term $|\mathbb{P}(e\in X_{T}^{v})-\pi_{G,v,\ell}^{-}(1_{e})|$ decays as $n\mathrm{e}^{-C_{\mathrm{d}}T_{v}/3}$ where $T_{v}$ is the number of updates inside the ball $B_{\ell}(v)$ by time $t$, which ultimately gives the claimed bound $\frac{1}{100m}$
Combining the above and going back to (E.1), we obtain
@@
In the remainder of the proof, we denote the $T$ established above for $\varepsilon=1/4$ by $T_{1/4}$. To get the claimed TV-distance bound for all $\varepsilon\geq\mathrm{e}^{-\Omega(n)}$, we modify the standard probability amplification argument using monotonicity. Formally, we prove the following inequality for all integer $k\geq 0$, whenever $T^{\prime}\geq kT_{1/4}$ (note that the second inequality follows from induction and the definition of $T_{1/4}$).
$\mathbb{P}(X_{kT_{1/4}}\neq\hat{X}_{kT_{1/4}}\mid{\mathcal{E}}_{<T^{\prime}})$ $\leq\frac{\mathbb{P}(X_{(k-1)T_{1/4}}\neq\hat{X}_{(k-1)T_{1/4}}\mid{\mathcal{E}}_{<T^{\prime}})\mathbb{P}(X_{T_{1/4}}\neq\hat{X}_{T_{1/4}})}{\mathbb{P}({\mathcal{E}}_{<T^{\prime}})}$ (E.2)
$\leq[4\mathbb{P}({\mathcal{E}}_{<T^{\prime}})]^{-k}$ (E.3)
@@
Thus, conditionally on ${\mathcal{E}}_{<T^{\prime}}$, it must be the case that both $X_{t}$, $\hat{X}_{t}$ has not coupled by time $(k-1)T_{1/4}$ (if they have, by the conditioning they would not disagree before time $T^{\prime}$), and also that $\hat{X}_{(k-1)T_{1/4}+t}$, $X^{\prime}_{t}$ has not coupled by time $t=T_{1/4}$, hence l.h.s. of (E.2) is at most

### v2

@@ Appendix C Remaining Proofs for Section 4
Write $T=\{i_{1},\dots,i_{k}\}$, and let $1\leq j\leq k$. Suppose we conditioned on the “outcome” for $R$-neighbourhoods of $i_{1},\dots,i_{j-1}$: that is, the number of surviving leaves, and which of the vertices in $\bigcup_{\ell<j}V_{i_{\ell}}$ belong to the largest component. Regardless of the previous neighbourhoods, there are at least $\chi n^{\prime}/2-d^{R}|S^{\prime}|\geq\chi n/3$ undetermined vertices in the largest component, and these are chosen, by symmetry, uniformly at random from the set of at most $n$ not yet determined vertices.
Now consider the number of surviving leaves, $L_{i_{j}}$, in the $R$-neighbourhood of $i_{j}$. By Corollary D.4, the probability that $L_{i_{j}}<Kd$ is $(\varphi p+1-p)^{d_{i_{j}}}+o_{R}(1)$ (as $(\varphi p+1-p)^{d_{i_{j}}}$ is the survival probability for the corresponding branching process). Taking $R$ large enough, we can get $o_{R}(1)\leq\epsilon_{1}/2$. Then, conditional on $L_{i_{j}}\geq Kd$, $|V_{i_{j}}|\geq K(d-1)$, and the probability that neither of these K(d1)K(d-1) belong to a u.a.r. subset of <n<n vertices of size $\geq\chi n/3$ is at most $(1-\chi/3)^{K(d-1)}\rightarrow 0$ as we increase $K$. Thus for $K$ large enough, this probability is at most $\epsilon_{1}/2$.
@@
We showed $\mathbb{P}(|T|<(1-\epsilon_{0})|S^{\prime}|)\leq n^{-C|S^{\prime}|}$, and $\mathbb{P}({\mathcal{C}}_{good}\mid|T|\geq(1-\epsilon_{0})|S^{\prime}|)\leq\mathrm{e}^{-\Omega(n)}$. (D.1) gives $\mathbb{P}(S^{\prime}\cap{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})=\emptyset\mid|T|\geq(1-\epsilon_{0})|S^{\prime}|,T)\leq\prod_{i\in T}[(\varphi p+1-p)^{d_{i}}+\epsilon_{1}]$.
It remains to show that, for $\epsilon_{0},\epsilon_{1}$ sufficiently small and $|T|\geq(1-\epsilon_{0})|S^{\prime}|$, we have that
@@
As a first step towards (D.2), note that for each $i\in S^{\prime}$, $d_{i}\leq d$, and since also $(\varphi p+1-p)<1$, it follows that $(\varphi p+1-p)^{d_{i}}+\epsilon_{1}\leq(\varphi p+1-p)^{d_{i}}(1+\epsilon_{1}/(\varphi p+1-p)^{d})$. Also, we can bound, for all sufficiently large $T$,
@@
Combining, we therefore get that
$\prod_{i\in T}[(\varphi p+1-p)^{d_{i}}+\epsilon_{1}]\leq[(1+\tfrac{\epsilon_{1}}{(\varphi p+1-p)^{d}})(\varphi p+1-p)^{d\epsilon_{0}}]^{|S^{\prime}|}\prod_{i\in S^{\prime}}(\varphi p+1-p)^{d_{i}}$
Note that by taking $\epsilon_{0},\epsilon_{1}\rightarrow 0$, we get $(1+\frac{\epsilon_{1}}{(\varphi p+1-p)^{d}})(\varphi p+1-p)^{d\epsilon_{0}}\rightarrow 1$, so in particular we can have $\epsilon_{0},\epsilon_{1}>0$ small enough so that expression becomes at most $1+\epsilon/2$, yielding (D.2).
@@
$\epsilon(\varphi p+1-p)^{d}/2]^{|S^{\prime}|}+(1+\epsilon/2)^{|S^{\prime}|}\prod_{i\in S^{\prime}}(\varphi p+1-p)^{d_{i}}\leq(1+\epsilon)^{|S^{\prime}|}\prod_{i\in S^{\prime}}(\varphi p+1-p)^{d_{i}},$
@@
Appendix E Remaining Proofs for Section 5
E.1. Proof of Lemma 5.3
In this section, we give the proofs of Equations (5.3) and (5.4).
Recall that $\nu_{\mathrm{o}}$ and $\rho_{\mathrm{o}}$ are respectively the vertex colour and the half-edge colour statistic for the ordered phase. Then let $P$ is an arbitrary sequence of vertices $(v_{0},v_{1},\dots,v_{\ell})$ with $q$-colouring $\tau=(\tau_{0},\dots,\tau_{\ell}$ and $\xi$ is a set of paired half-edges that turn $v_{0},\dots,v_{\ell}$ into a path in that order. Then ${\mathcal{E}}_{P,\tau,\epsilon}$ is the event that $v_{0},v_{1},\dots,v_{\ell}$ form a path in $\hat{\boldsymbol{G}}$ using the half-edges prescribed by $\epsilon$ and that their colours are as prescribed by $\tau$. We claim the following asymptotic bound on the probability of ${\mathcal{E}}_{P,\tau,\xi}$:
Proof of Equation (5.3) .
Recall that the graph $\hat{\boldsymbol{G}}$ has $n_{i}=n\nu_{i}$ vertices of colour $i\in[q]$ and the number of edges between colour classes $i$ and $j$ with $i\neq j$ is given by $2m_{ij}:=dn\rho_{ij}$ if $i\neq j$, and $m_{ii}:=\frac{dn}{2}\rho_{ii}$ if $i=j$, cf. (2.4); we further have that $\nu$ and $\rho$ satisfy $\|\nu-\nu_{\mathrm{o}}\|_{1}+\|\rho-\rho_{\mathrm{o}}\|_{1}=O(n^{-1/3})$. The graph $\hat{\boldsymbol{G}}$ is uniformly distributed conditioned on these statistics. In particular, the number of ways to match the half-edges consistent with the vertex statistics $\{n_{i}\}$ and edge-statistics $\{m_{ij}\}$ is given by
We have
$\frac{\mathcal{N}^{\prime}}{\mathcal{N}}=\frac{(n-(\ell+1))!}{n!}\prod_{i\in[q]}\frac{n_{i}!}{(n_{i}-\ell_{i})!}\,\prod_{i\in[q]}\frac{(M_{i})!}{(dn_{i})!}\,\prod_{i\in[q]}\frac{2^{m_{ii}}(m_{ii})!}{2^{m_{ii}^{\prime}}(m_{ii}^{\prime})!}\,\prod_{i<j}\frac{(2m_{ij})!}{(2m_{ij}^{\prime})!}.$ (E.1)
It is a matter of a few manipulations now to massage the last expression into (5.3). ∎
We next prove Equation (5.4), i.e.
[…]
where $\ell_{1}(\tau)$ is the number of vertices with colour 1 under $\tau$ and $\hat{\varphi}_{1}$ is the extinction probability of the modified Galton-Waston process (where the root has degree d2d-2, non-root vertices have degree $d$), as defined in (5.1).
Proof of Equation (5.4).
Condition on ${\mathcal{E}}_{P,\tau,\xi}$ and on the set of edges $E_{v}$ in $\hat{\boldsymbol{G}}$ incident to $v$. Let $\mathbf{d}$ be the degree sequence of the graph $\hat{\boldsymbol{G}}\backslash v=(V\backslash v,E\backslash(E_{P}\cup E_{v}))$. For a colour $i\in[q]$, let $\boldsymbol{V}_{\!i}$ denote the $i$-th colour class of $\v\hat{\boldsymbol{G}}\backslash$ v under $\hat{\boldsymbol{\sigma}}_{\mathrm{o}}$, $\mathbf{d}_{i}$ be the subsequence of $\mathbf{d}$ corresponding to vertices in $\boldsymbol{V}_{\!i}$, $\boldsymbol{E}_{i}$ be the set of monochromatic edges within $\boldsymbol{V}_{\!i}$, and $\boldsymbol{F}_{i}=\boldsymbol{F}\cap\boldsymbol{E}_{i}$. Conditioned on $|\boldsymbol{F}_{i}|=f_{i}$, the graph $\boldsymbol{G}_{i}=\big(\boldsymbol{V}_{\!i},\boldsymbol{F}_{i}\backslash(E_{P}\cup E_{v})\big)$ has the same distribution as the exact-edge model $\tilde{\boldsymbol{G}}_{n_{i},\mathbf{d}_{i},f_{i}}$ where $n_{i}=|\boldsymbol{V}_{\!i}|=n\nu_{\mathrm{o},i}+o(n)$. Note that for all $i\in[q]$, there are at most $\ell+1+d=O(n^{1/6})$ vertices in $\boldsymbol{d}_{i}$ with degree $\leq d$ (all other vertices have degree $d$).
Since $|\boldsymbol{E}_{i}|=\frac{dn}{2}\rho_{\mathrm{o},ii}+o(n)$, by a Chernoff bound we have that $|\boldsymbol{F}_{i}|=p\frac{dn}{2}\rho_{\mathrm{o},ii}+o(n)$ with probability $1-\mathrm{e}^{-\Omega(n)}$. It follows from Lemma 5.2, applied with densities $p_{i}=p\frac{\rho_{\mathrm{o},ii}}{\nu_{\mathrm{o},i}}$ for $i\in[q]$, that ${\boldsymbol{G}}_{1}$ is in the supercritical regime while ${\boldsymbol{G}}_{i}$ for $i\in\{2,\ldots,q\}$ is in the subcritical regime. Hence, it follows from Lemma 5.1 that w.p. $1-\mathrm{e}^{-\Omega(n)}$, the largest component $\mathcal{C}$} in $\boldsymbol{G}_{1}$ has size $\Omega(n)$ and the remaining components have all size $O(n^{2/3})$. Moreover, from Lemma 5.1 part (2) we obtain that [^4] Note that, because the edges in $E_{v}$ are excluded from $\hat{\boldsymbol{G}}\setminus v$, all but at most $d+2$ vertices of the path $P$ have degree d2d-2 in $\hat{\boldsymbol{G}}\setminus v$; the contribution of vertices with degree <d2<d-2 is at most a constant which is absorbed into the $(1+\frac{\varepsilon}{3})^{\ell}$ factor in (E.2).
$\mathbb{P}[\,{\mathcal{C}}\mbox{ avoids the path }P\mid{\mathcal{E}}_{P,\tau,\xi},E_{v}]\leq(1+\tfrac{\varepsilon}{3})^{\ell}(p\varphi+1-\varphi)^{(d-2)\ell_{1}}\leq(1+\tfrac{\varepsilon}{3})^{\ell+1}\hat{\varphi}_{1}^{\ell_{1}}.$ (E.2)
[…]
Moreover, the length of the path $P$ is $\leq n^{1/6}$, so $|{\mathcal{C}}^{\prime}|-|{\mathcal{C}}|=O(n^{5/6})$. Similarly, any component in $\hat{\boldsymbol{G}}_{v}(\boldsymbol{F})$ other than $\mathcal{C}$$^{\prime}$ has size $O(n^{2/3})$, and in particular $\mathcal{C}$$^{\prime}$ is the largest component in $\hat{\boldsymbol{G}}_{v}(\boldsymbol{F})$. Combining this with (E.2) and (E.3), we obtain (5.4). ∎
E.2. Remaining proofs for Lemma 5.4
Next we prove Lemma 5.4 which we restate for convenience.
Lemma 5.4.
Let $d\geq 3$, $q\geq(5d)^{5}$ and $\beta\geq\beta_{c}$. Then, for all $i\in[q]$, it holds that ${(1-\sqrt{\hat{\varphi}_{1}})}\frac{\rho_{\mathrm{o},i1}}{\nu_{\mathrm{o},i}}>1-\frac{1}{(d-1)^{5}}$.
Proof.
$\frac{\rho_{\mathrm{o},{i1}}}{\nu_{\mathrm{o},i}}=1-R\mbox{ where }R:=\tfrac{t^{d+1}+(q-2)t^{d}-(q-1)t}{(t^{d}-t)(t^{d-1}+q-1)}$ (E.4)
To upper bound $\hat{\varphi}_{1}$, we need first to find the solution of $\varphi_{1}=(p_{1}\varphi_{1}+1-p_{1})^{d-1}$ for $p_{1}=p\frac{\rho_{\mathrm{o},11}}{\nu_{\mathrm{o},1}}$. Similarly to above, we have $\rho_{\mathrm{o},11}=\frac{\mathrm{e}^{\beta}t^{2(d-1)}}{(t^{d-1}+q-1)^{2}+(\mathrm{e}^{\beta}-1)(t^{2d-2}+(q-1))}$, so using the expression for $\mathrm{e}^{\beta}-1$ from (2.1) we obtain after some algebra that $p_{1}=\frac{t^{d-1}(t-1)}{t^{d}-t}$ and $\varphi_{1}=1/t^{d-1}$. Since $t\geq t_{c}$ for $\beta\geq\beta_{c}$, we obtain that $\varphi_{1}\leq\frac{1}{(q-1)^{2(d-1)/d}}$. It follows that $\hat{\varphi}_{1}=(p_{1}\varphi_{1}+1-p_{1})^{d-2}=\varphi_{1}^{(d-2)/(d-1)}\leq\frac{1}{(q-1)^{2(d-2)/d}}$, so $\sqrt{\hat{\varphi}_{1}}\leq\frac{1}{(q-1)^{(d-2)/d}}$. For $q\geq(5d)^{5}$, numerical calculations give that this is less than $\frac{1}{2(d-1)^{5}}$ for all $d\geq 3$ (the inequality is tighter for small values of $d$). A similar argument shows that $R<\frac{1}{2(d-1)^{5}}$ for all $d\geq 3$ when $q\geq(5d)^{5}$; to see this, we have from the expression in (E.4) that $R\leq R^{\prime}$ where $R^{\prime}:=\tfrac{t+q-1}{t^{d-1}+q-1}$. We have that $R^{\prime}$ is decreasing in $t$, so using that $t\geq t_{c}$, we obtain that $R^{\prime}\leq\tfrac{t_{c}+q-1}{t^{d-1}_{c}+q-1}=\frac{1}{(q-1)^{(d-2)/d}}$; from the same numerical inequality as above, we therefore conclude that $R$ is less than $\frac{1}{2(d-1)^{5}}$ for all $d\geq 3$ when $q\geq(5d)^{5}$. Hence, $\frac{1}{(d-1)^{5}}>\sqrt{\hat{\varphi}_{1}}+R$, as needed. ∎
E.3. Remaining Proofs for Section 5.1
Finally, we give the proofs of Lemmas 5.6 and 5.7, which we restate for convenience.
Lemma 5.6.
$\big|\pi_{G,v,\ell}^{+}(1_{e})-\pi_{G,\mathrm{o}}(1_{e})\big|\leq\pi_{G,\mathrm{o}}\big(\mbox{\emph{no wired boundary around $v$ in $B_{\ell}(v)$}}\big)+\mathrm{e}^{-\Omega(n)}.$
Proof.
Let $\boldsymbol{F}_{\mathrm{o}}\sim\pi_{G,\mathrm{o}}$ and $\boldsymbol{F}^{+}\sim\pi_{G,v,\ell}^{+}$.
Let $\mathcal{O}$ be the event that $\boldsymbol{F}_{\mathrm{o}}$ contains at least $n\mathfrak{m}_{\mathrm{o}}(\beta_{c})-n\rho/4$ edges outside of the ball (where $\rho=\mathfrak{m}_{\mathrm{o}}(\beta_{c})-\mathfrak{m}_{\mathrm{d}}(\beta_{c})$ as in Definition 3.1). Since $|E(B_{\ell}(v))|=O(n^{1/2})$, for $n$ large enough we have $\mathbb{P}(\overline{\mathcal{O}})\leq\mathbb{P}(|\boldsymbol{F}_{\mathrm{o}}|\leq n\mathfrak{m}_{\mathrm{o}}-n\rho/8)$, which is by Corollary B.2, whp over $G$, bounded by $\mathrm{e}^{-\Omega(n)}$. Note that conditional on $\mathcal{O}$, regardless of the state of the edges in the ball $B_{\ell}(v)$, the resulting configuration is ordered and thus we can thus drop the conditioning on the ordered phase.
By the monotonicity property of the random cluster model (see e.g. [18, Appendix B]), we get that $\pi_{G,v,\ell}^{+}$ stochastically dominates $\pi_{G,\mathrm{o}}(\cdot\mid\mathcal{O})=\pi_{G}(\cdot\mid\mathcal{O})$ on $B_{\ell}(v)$, hence conditional on $\mathcal{O}$ we can perfectly couple $\boldsymbol{F}_{\mathrm{o}}\sim\pi_{G,\mathrm{o}}$ and $\boldsymbol{F}^{+}\sim\pi_{G,v,\ell}^{+}$ such that $\boldsymbol{F}_{\mathrm{o}}\cap E(B_{\ell}(v))\subseteq\boldsymbol{F}^{+}\cap E(B_{\ell}(v))$. Then we can bound the probability that $\boldsymbol{F}_{\mathrm{o}}$ and $\boldsymbol{F}^{+}$ disagree on $e$ by
Conditional on ${\mathcal{E}}_{\boldsymbol{F}_{\mathrm{o}},\mathcal{S}},\mathcal{O}$, we can sample $\boldsymbol{F}_{\mathrm{o}}$, $\boldsymbol{F}^{+}$ by first generating the edges outside of the component of $v$ in $G\setminus\mathcal{S}$ – call these edges $E_{O}$ – and then, conditional on $E_{O}$ and ${\mathcal{E}}_{\boldsymbol{F}_{\mathrm{o}},\mathcal{S}}$ the remainder of the edges. Note that for $\mathcal{S}\in\mathcal{S}_{B}$, the events $\mathcal{S}$ is a wired boundary in $\boldsymbol{F}_{\mathrm{o}}$” (which implies it is a wired boundary also in $\boldsymbol{F}^{+}$) and $\mathcal{O}$ are determined by $\boldsymbol{F}_{\mathrm{o}}\cap E_{O}$.
Thus, conditional on ${\mathcal{E}}_{\boldsymbol{F}_{\mathrm{o}},\mathcal{S}}$, $\mathcal{O}$, and then any (possible) subconfiguration of $(\boldsymbol{F}_{\mathrm{o}},\boldsymbol{F}^{+})$ on $E_{O}$, they both have a wired boundary $\mathcal{S}$. Then by [18, Observation 51], the marginals edges in $E\setminus E_{O}$ agree, hence since $e\not\in E_{O}$ and by the fact that $(\boldsymbol{F}_{\mathrm{o}},\boldsymbol{F}_{+})$ were perfectly coupled, the probability of disagreeing on $e$ conditional on ${\mathcal{E}}_{\boldsymbol{F}_{\mathrm{o}},\mathcal{S}}$ and $\mathcal{O}$ is zero. ∎
Lemma 5.7.
Let $G=(V,E)$ be a graph, let $v\in V$ be an arbitrary vertex, and let $\ell\geq 1$ an integer. For any $F\subseteq E$, let $F_{v}$ be the set of edges in $F$ that are incident to $v$ and let $G_{v,F}:=(V\setminus\{v\},F\setminus F_{v})$. Suppose that all simple length-$\ell$ paths from $v$ in $G$ intersect the largest component of $G_{v,F}$. Then $F$ has a wired boundary around $v$ in $B_{\ell}(v)$.
Proof of Lemma 5.7.
Consider the set of paths $\mathcal{P}$ of length $\ell$ in $G$ that start in $v$.
Claim 2. No vertices of ${\mathcal{C}}_{1}$ belong to ${\mathcal{C}}_{v}$.
Appendix F Proof of Theorems 3.4 and 3.5
@@ Appendix E Remaining Proofs for Section 5
In this section, we prove Theorems 3.4 and 3.5. The proofs of these follow closely those in [18] (which are in turn similar to previous arguments in [20, 9]), so we outline the key ideas and only included detail on the probability amplification (which is not included therein).
@@
Also, for $\ell$ as above, there exists a constant $L>0$ such that for every vertex $v$, $B_{\ell}(v)$ is $L$-treelike, see, e.g., [9, Fact 2.3.]. Consider any graph $G$ that satisfies these properties.
We next outline the theorem for $\varepsilon=\tfrac{1}{4}$, the full details for this part can be found in [18]. Then, we prove that probability amplification applies for any $\varepsilon=\mathrm{e}^{-\Omega(n)}$. Consider two copies of the random-cluster dynamics: (1) $(X_{t})_{t\geq 0}$ starting from all-out, and (2) $(\hat{X}_{t})_{t\geq 0}$ starting from $\pi_{G,\mathrm{d}}$ which is restricted to the disordered phase by ignoring updates that would make the configuration not disordered. Note that the stationary distribution of $\hat{X}_{t}$ is $\pi_{G,\mathrm{d}}$, and thus for all $t\geq 0$, $\hat{X}_{t}\sim\pi_{G,\mathrm{d}}$.
@@
$\mathbb{P}(X_{T}\neq\hat{X}_{T})\leq$ $m\mathbb{P}(\overline{{\mathcal{E}}_{<T}})+\sum_{e\in E(G)}\mathbb{P}(\mathds{1}\{e\in X_{T}\}\neq\mathds{1}\{e\in\hat{X}_{T}\}\mid{\mathcal{E}}_{<T}).$ (F.1)
@@
where $\pi^{-}_{G,v,\ell}$ is the distribution on $B_{\ell}(v)$ conditioned on an all-out boundary. Since $\mathbb{P}(e\in\hat{X}_{T})=\pi_{G,\mathrm{d}}(1_{e})$, WSM within the disordered phase bounds the third term by $\frac{1}{100m}$. The second term is also bounded by $\frac{1}{100m}$ using mixing time estimates, in this case log-Sobolev inequalities. We don’t give the full details here (which, as mentioned before, can also be found in [18, 20, 9]) but the idea is that, since the neighbourhood of $v$ is treelike (up to removing a constant number of edges), the log-Sobolev constant for $\pi_{G,v,\ell}^{-}(1_{e})$ is a constant factor away from that on the regular tree. This in turn is just percolation so the log-Sobolev constant for the chain $X_{v}^{t}$ is at most $C_{\mathrm{d}}/|V(B_{\ell}(v))|$ (in the ordered regime, the bound for the log-Sobolev constant on the tree with + boundary condition is given in [7]). Since $\pi_{G,v,\ell}^{-}$ is the stationary distribution of the chain $X_{v}^{t}$, the term $|\mathbb{P}(e\in X_{T}^{v})-\pi_{G,v,\ell}^{-}(1_{e})|$ decays as $n\mathrm{e}^{-C_{\mathrm{d}}T_{v}/3}$ where $T_{v}$ is the number of updates inside the ball $B_{\ell}(v)$ by time $t$, which ultimately gives the claimed bound $\frac{1}{100m}$.
Combining the above and going back to (F.1), we obtain
@@
In the remainder of the proof, we denote the $T$ established above for $\varepsilon=1/4$ by $T_{1/4}$. To get the claimed TV-distance bound for all $\varepsilon\geq\mathrm{e}^{-\Omega(n)}$, we modify the standard probability amplification argument using monotonicity. Formally, we prove the following inequality for all integer $k\geq 0$, whenever $T^{\prime}\geq kT_{1/4}$ (the second inequality follows from induction and the definition of $T_{1/4}$).
$\mathbb{P}(X_{kT_{1/4}}\neq\hat{X}_{kT_{1/4}}\mid{\mathcal{E}}_{<T^{\prime}})$ $\leq\frac{\mathbb{P}(X_{(k-1)T_{1/4}}\neq\hat{X}_{(k-1)T_{1/4}}\mid{\mathcal{E}}_{<T^{\prime}})\mathbb{P}(X_{T_{1/4}}\neq\hat{X}_{T_{1/4}})}{\mathbb{P}({\mathcal{E}}_{<T^{\prime}})}$ (F.2)
$\leq[4\mathbb{P}({\mathcal{E}}_{<T^{\prime}})]^{-k}$ (F.3)
@@
Thus, conditionally on ${\mathcal{E}}_{<T^{\prime}}$, it must be the case that both $X_{t}$, $\hat{X}_{t}$ has not coupled by time $(k-1)T_{1/4}$ (if they have, by the conditioning they would not disagree before time $T^{\prime}$), and also that $\hat{X}_{(k-1)T_{1/4}+t}$, $X^{\prime}_{t}$ has not coupled by time $t=T_{1/4}$, hence l.h.s. of (F.2) is at most